Grade 12th passGeneral Physics

What is the de -Broglie wavelength associated with an alpha particle accelerated by a potential difference of 25KV?

Anand G
8 Years agoGrade 12th pass

1 Answer

vivek jaiswal
8 Years ago
Wavelength=h/√2mqvh=6.63×10^-34jsV(given)=25000vq=2p=2×1.6×10^-19cm=4amu=4×1.6×10^-27Putting these valueWavelength=6.5×10^-14m